If 𝑓(π‘₯) is a differentiable and 𝑔(π‘₯) = π‘₯ 𝑓(π‘₯) use the definition of the derivative to show



that 𝑔′(π‘₯) = π‘₯𝑓'(π‘₯) + 𝑓(π‘₯).

1
Expert's answer
2022-03-28T07:05:21-0400

Given a function "f(x)" , the derivative (by definition) is:


"f^{\prime}(x)=\lim _{h \rightarrow 0} \frac{f(x+h)-f(x)}{h}"

Now , given "\ud835\udc54(\ud835\udc65) = \ud835\udc65 \ud835\udc53(\ud835\udc65)" ,


"\begin{aligned}
g'(x) = & \lim_{h \rightarrow 0} \frac{(x+h)f(x+h) - xf(x)}{h}\\
=& \lim_{h \rightarrow 0} \frac{xf(x+h)+hf(x+h) - xf(x)}{h}\\
=& \lim_{h \rightarrow 0} \frac{xf(x+h)- xf(x)+hf(x+h)}{h}\\
=& \lim_{h \rightarrow 0} \frac{x(f(x+h)- f(x))+hf(x+h)}{h}\\
=& \lim_{h \rightarrow 0} \frac{x(f(x+h)- f(x))}{h} + \lim_{h \rightarrow 0} \frac{hf(x+h)}{h}\\
=& \lim_{h \rightarrow 0} x\Bigg(\frac{f(x+h)- f(x)}{h}\Bigg) + \lim_{ h\rightarrow 0} f(x+h)\\
g'(x)= & xf'(x) + f(x)
\end{aligned}"

Q.E.D